OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos x = sin x

OpenStudy (anonymous):

pi/4

OpenStudy (ghazi):

just use tan x= 1 and solve for x

OpenStudy (anonymous):

x has to be pi on 4 for both to be equal

OpenStudy (anonymous):

pi/4 and 7pi/4 right?

OpenStudy (anonymous):

u cant have 7pi on 4

OpenStudy (anonymous):

cause the 4th quadrant is positive for cos values but negative for sin values

OpenStudy (anonymous):

5pi/4

OpenStudy (anonymous):

ucant...

OpenStudy (anonymous):

ou can rewrite sin2x using the double angle formula to get: cosx = 2*sinx*cosx... subtract cosx on both sides 2sinxcosx - cosx = 0... factor cosx(2sinx - 1) = 0 Set each factor equal to zero to get two sets of roots: cosx = 0 ==> x = {π/2,3π/2} 2sinx - 1 = 0 ==> sinx = 1/2 ==> x = {π/6, 5π/6}

OpenStudy (anonymous):

ur interval is form 0 < x < 2pi

OpenStudy (anonymous):

the only x values that is common to both cos x = sin x is when x = pi/4

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

3pi/4, 5pi/4 pi/4,7pi/4 3pi/4, 7pi/2 pi/4, 5pi/4

OpenStudy (anonymous):